# A Sparse Graph Formulation for Efficient Spectral Image Segmentation

### v1

@@ Abstract
Spectral Clustering is one of the most traditional methods to solve segmentation problems. Based on Normalized Cuts, it aims at partitioning an image using an objective function defined by a graph. Despite their mathematical attractiveness, spectral approaches are traditionally neglected by the scientific community due to their practical issues and underperformance. In this paper, we adopt a sparse graph formulation based on the inclusion of extra nodes to a simple grid graph. While the grid encodes the pixel spatial disposition, the extra nodes account for the pixel color data. Applying the original Normalized Cuts algorithm to this graph leads to a simple and scalable method for spectral image segmentation, with an interpretable solution. Our experiments also demonstrate that our proposed methodology over performs traditional spectral algorithms for segmentation.
@@ 1 Introduction
Image Segmentation refers to the problem of grouping image pixels in meaningful regions with the goal of understanding or summarizing the visual content in that image [1]. Because of that, image segmentation has found applications in many domains, such as medical image analysis, autonomous driving, and photo editing to name a few [1]. This intense industrial appeal led to the development of many algorithms, which tackled the problem using variational [2], graphical [3, 4, 5], statistical [6] and, more recently, deep leaning techniques [7].
Spectral clustering [8], pioneered by Normalized Cuts [3], provides an algorithm for image segmentation that stands out for its methodological consistency backed by graph theory and computational linear algebra. It consists of encoding the image data into a graph then using spectral techniques to approximately find the cut corresponding to the final segmentation. While popular in its early years, the initial formulation of Normalized Cuts had that would hinder its broad adoption by the scientific community. The issues found were mostly (1) computational, as it required the construction and decomposition of large matrices [9], and (2) practical, as it performed worse than other algorithms [10].
Many strategies were approached to overcome the issues [9, 11, 12, 13]. In this paper, we approach them by adopting a sparse graph consisting of a grid graph that encodes the pixel spatial locations and of a few extra nodes corresponding to the pixel intensities [2, 4]. We show that the Normalized Cut solution found using this graph encourages pixel groupings that linearly balances spatial coherence (neighbors generally belong to the same segment) and color cohesion (pixels in a segment have colors in the same subset of available colors) [14]. Furthermore, we also demonstrate that our proposed spectral method is scalable and easily tunable, besides being easy to implement. Finally, we compare our method to other traditional spectral techniques for segmentation and observe that it outperforms them in terms of runtime and quality.
The next sections are organized as follows: In Section 2, we show an overview of Normalized Cuts, its main algorithm, and drawbacks. In Section 3, we present our proposed graph construction and characterize the solution of Normalized Cuts when optimized to it. Experimental results are presented in Section 4 along with a brief discussion on them. We finish with conclusions and future work in Section 5.
@@ 2.1 Normalized Cuts
Let $I$ be an image of $n$ pixels and $k$ unique colors. Let $G=(V,E,w)$ be an undirected weighted graph where each node corresponds to a pixel and the weight function $w(i,j)$ evaluates the similarity between pixels $i$ and $j$. A binary segmentation of $I$ is achieved by finding a partition $(A,B)$ of $V$ that minimizes the cut value $\operatorname{\operatorname{Cut}}(A,B)=\sum_{i\in A,j\in B}w(i,j)$.
This approach however favors unbalanced partitions, leading to undesirable segmentations. One successful way of addressing this issue is to consider the normalized cut value $\operatorname{\operatorname{NCut}}(A,B)$ [3] instead,
$\operatorname{\operatorname{NCut}}(A,B)=\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(A)}+\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(B)}.$
Here $\operatorname{\operatorname{Vol}}(A)$ is a measure of the “volume” of $A$ defined as $\operatorname{\operatorname{Vol}}(A)=\sum_{i\in A,j\in V}w(i,j)$. Later, we use $\operatorname{\operatorname{Vol}}(V)=\operatorname{\operatorname{Vol}}(A)+\operatorname{\operatorname{Vol}}(B)$ to represent the volume of all edges in $G$.
Let $W$ be the weighted adjacency matrix of $G$ and $D$ be the diagonal degree matrix with $D(i,i)=\sum_{j\in V}w(i,j)$. The Laplacian matrix of $G$ is then defined as $L=D-W$. Following the derivation presented in [15], we can write the problem of minimizing the $\operatorname{\operatorname{NCut}}$ of $G$ as:
$\min_{\mathbf{x}}\quad$ $\mathcal{E}(\mathbf{x}|G),$ (1)
@@
2.2 Classical Edge Weight for Image Segmentation
A crucial step in developing efficient segmentation solvers using $\operatorname{\operatorname{NCut}}$ is to define how the function $w$ is computed. In the original $\operatorname{\operatorname{NCut}}$ work and in early subsequent work [16, 17, 18], the $w$ combined two grouping cues in a single value:
@@ 2.2 Classical Edge Weight for Image Segmentation
where $I(j)$ and $X(j)$ are the appearance (such as the color) of pixel $j$ and its spatial location on image $I$, respectively, $\mathds{1}(\cdot)$ is the indicator function, and $r$, $\sigma_{I}$ and $\sigma_{X}$ have fixed values.
2.3 Drawbacks of the Traditional Approach
The $\operatorname{\operatorname{NCut}}$ formulation based on the above weight function presents several drawbacks. For one, it is computationally inefficient to construct the above graph for high resolution images, since it requires visiting each pixel and their neighboring locations withing a radius $r$. Secondly, this graph is potentially dense, which poses memory and runtime difficulties during eigenvector decomposition. $\operatorname{\operatorname{NCut}}$ using the weight as in Eq. 4 is also known to be inefficient, even for small images [10].
Moreover, calibrating the values for $\sigma_{X}$ and $\sigma_{I}$ is challenging [14]. Finally, the optimization in Eq. 1 on the graph constructed using Eq. 4 lacks interpretability [14]. The only understanding we can extract from its solution is that it encourages similar pixels to be grouped together, which is commonplace for any spectral clustering algorithm.
@@ 3.1 New Graph Formulation
Inspired by [4] and [2], we show that a sparse graph formulation efficiently tackles the issues with the classical $\operatorname{\operatorname{NCut}}$ formulation. Our graph construction starts by considering a grid graph $G_{\text{grid}}$ whose vertices correspond to image pixels, with edges of weight $\mu>0$ between each pair of neighboring pixels. From $G_{\text{grid}}$, we create a new graph $G_{\text{all}}$ by adding $k$ extra nodes, each corresponding to one of the $k$ unique colors in $I$. Each node on $G_{\text{grid}}$ is then connected to the extra node corresponding to its color with an edge of unitary weight. This graph formulation is computationally attractive for its sparsity. Assuming $k\ll n$, $G_{\text{all}}$ only requires $O(n)$ edges. When segmenting $I$, we only make use of the assignment vector for the grid nodes.
@@ 3.2 Interpretation of the Cut on $G_{\text{all}}$
Table 1: Quantitative results on synthetic images in terms of JAC and runtime. The data was generated using four different noise models on three different ground truth patterns. The lower the JAC value, the better.
[…]
tab_synt
Original
Proposed
NCut
MixNCut
MGSC
Fig. 1: Qualitative results on synthetic images. Top: Gaussian noise, $\sigma_{g}^{2}=1.5$. Bottom: Salt & Pepper, $d=0.7$.
@@
Now, because of the grid nature of $G_{\text{grid}}$, the second term in Eq. 6 accounts for the number of neighboring pixel pairs that are assigned to different segments:
@@
Minimizing $\mathcal{E}(\mathbf{z}|G_{\text{all}})$ for $\mathbf{x}$ and $\mathbf{y}$, results in finding a partition of $G_{\text{all}}$ that trades off between spacial coherence (neighboring pixels should belong to the same segment) and color cohesion (pixels in a segment should have colors belonging to the same subset of available intensities) via the value of $\mu$. Minimizing the multiplicative factor in Eq. 9 also encourages a balanced partition of the nodes in $G_{\text{all}}$.
(a) Image
(b) $\mu=5$ ($8.63s$)
(c) $\mu=50$ ($23.22s$)
(d) $\mu=300$ ($104.56s$)
Fig. 2: Qualitative results on RBG images for various $\mu$. (b)-(d) show the eigenvector from Eq. 3 (left) and the corresponding segmentation (right). The image sizes are $800\times 618$ (top) and $800\times 524$ (bottom). Average runtimes are shown in parentheses.
3.3 Improvements
3.3.1 Soft color linkage
Despite the simplicity and computational efficiency of the above result, it does not explicitly use the pixel color intensities. This means that we would still achieve the same solution to Eq. 1 even if we randomly relabeled the set of available colors. In our experiments, we test the performance of a softer version of the graph described in the previous sections. In this new graph construction, each node $i$ in $G_{\text{grid}}$ is linked to all extra nodes with an edge of weight given by
$w^{\prime}(i,c)=\exp\left(\frac{||I(i)-c||^{2}}{2\sigma^{2}}\right)$ (10)
instead of the unitary weight from the original approach. The interpretation of the solution of Eq. 1 under this new graph is left for future work.
3.3.2 Case when $k$ is large
Our method assumes $k$, the number of unique colors in $I$, to be small compared to $n$. Although this assumption is satisfied for gray scale images, it becomes unrealistic for colored ones. In this case, we vectorize the RGB color space using $K$-means clustering. The weights in Eq. 10 are then computed using the average RGB color in each cluster. This approach also alleviates the issue posed in the previous section.
@@ 4.1 Experimental Setup
For performance evaluation, we assessed the segmentations via the Jaccard Dissimilarity measure (JAC), defined as:
$\text{JAC}=1-\max(J(R,S),J(R,\neg S))$ (11)
where $R$ and $S$ are the ground truth image and segmentation, respectively, and $J(R,S)=|R\cap S|/|R\cup S|$ is the Jaccard Index [19]. $\neg S$ is the segmentation with labels swapped.
@@
For our synthetic results, we contaminate three different background and foreground square patterns (intensities of either 0 and 1) of side length $\ell$ with either additive zero mean and variance $\sigma_{g}^{2}$ Gaussian noise or Salt & Pepper with noise density $d$. The generated images are then rescaled and discretized, so the color intensities are integers in $[0,255]$. We also test our method using two RGB images from the Alpha Matting dataset [20]. For each of them, we set $k=256$ in their preprocessing stage (ref. Section 3.3.2).
We compare our proposed method with $\operatorname{\operatorname{NCut}}$ and two more spectral methods for unsupervised segmentation: Multi Graph Spectral Clustering (MGSC) [21] and MixNCut [14]. For the later two, we create two separate graphs, a grid one for pixel locations and a fully connected one for color data, as described in [14]. When using all these techniques, we refrain from any sparsification algorithm for simplicity. For $\operatorname{\operatorname{NCut}}$, we also set $r=\infty$ in Eq. 4 for experimental simplicity and so that we can compare similar graph formulations.
@@
The experiments ran on an Apple M1 Pro Chip (8 core CPU, 14 core GPU) with 16 Gb of RAM memory and were coded in Python 3.9.7 using Numpy 1.21.4 and Scipy 1.7.3111A demo code for this paper is available at www.github.com/rahul-palnitkar/Sparse-Graph-Spectral-Segmentation..

### v3

@@ Abstract
Spectral Clustering is one of the most traditional methods for solving segmentation problems. Based on Normalized Cuts, it partitions an image using an objective function defined by a graph. Despite their mathematical attractiveness, spectral approaches have traditionally been neglected by the scientific community because of their practical issues and underperformance. In this paper, we adopt a sparse graph formulation based on the inclusion of extra nodes in a simple grid graph. While the grid encodes the spatial disposition of the pixels, the extra nodes account for the pixel color data. Applying the original Normalized Cuts algorithm to this graph leads to a simple and scalable method for spectral image segmentation, with a rich interpretable solution. Our experiments also show that our proposed methodology overperforms both traditional and modern unsupervised algorithms for segmentation in both real and synthetic data, establishing competitive results for many unsupervised segmentation datasets.
@@ 1 Introduction
Image Segmentation refers to the problem of grouping image pixels in meaningful regions with the goal of understanding or summarizing the visual content of that image [1]. Because of that, image segmentation has found applications in many domains, such as medical image analysis, autonomous driving, and photo editing, to name a few [1]. This intense industrial appeal led to the development of many algorithms that tackled the problem using variational [2], graphical [3, 4], statistical [5] and, more recently, deep learning techniques [6].
Normalized Cuts ($\operatorname{\operatorname{NCut}}$) [3] provides one of the most classical algorithms for image segmentation, and it stands out for its methodological consistency backed by graph theory and computational linear algebra. It consists of encoding the image data into a graph and then using spectral techniques to approximately find the cut corresponding to the final segmentation. It has been a lasting algorithm for vision problems [7] and is currently receiving prodigious research interest due to the emergence of Vision Transformers [8], due to its ability to automatically detect objects and segment images using Transformer-based image embeddings, such as those arising from self-distillation (DINO) approaches [9, 10, 11].
Despite its persistent theoretical appeal, $\operatorname{\operatorname{NCut}}$’s initial formulation rapidly met with drawbacks that would hinder its broader adoption by the scientific community, although many strategies were approached to overcome them [7, 12]. Historically, the issues were mainly (1) computational, as it required handling large matrices [7, 13], and (2) practical, as it performed worse than other unsupervised segmentation algorithms [14, 15]. Some of these issues were amended with the use of rich and low-resolution self-supervised transformer features [16] at the pixel level, followed by the standard $\operatorname{\operatorname{NCut}}$ approach [17, 11]. Despite that, these methods still scale poorly, only being shown to perform well on the low-resolution scale typical of features arising from vision transformers, often needing post-processing algorithms such as Conditional Random Fields (CRF) or Bilateral Solvers [11] for edge refinement. Finally, none of these approaches considers the interpretability of their optimization problem.
In this paper, we address all these issues by adopting a sparse graph consisting of a grid graph that encodes the pixel spatial locations and a few extra nodes corresponding to the pixel intensities [2]. We show that the $\operatorname{\operatorname{NCut}}$ solution found using this graph is interpretable, as it encourages pixel groupings that linearly balance spatial coherence (neighbors generally belong to the same group) and color cohesion (pixels in a segment have colors in the same subset of colors) [18]. We also demonstrate that this simple spectral formulation is scalable, and that its only parameter has a clear interpretation, which facilitates its tuning. Finally, we show that using our methodology to segment images based on their extracted DINO features outperforms recent approaches to use this data in the context of spectral segmentation [19, 11]. We compare our method with other traditional and modern unsupervised techniques for segmentation and observe that our simple methodology outperforms its spectral counterparts in terms of runtime and quality, while also being competitive against the other methods.
@@ 2.1 Normalized Cuts
Let $I$ be an image of $n$ pixels and $k$ unique colors. Let $G=(V,E,w)$ be an undirected weighted graph where each node corresponds to a pixel, and the weight function $w(i,j)$ evaluates the similarity between pixels $i$ and $j$. A binary segmentation of $I$ is achieved by finding a partition $(A,B)$ of $V$ that minimizes the cut value $\operatorname{\operatorname{Cut}}(A,B)=\sum_{i\in A,j\in B}w(i,j)$.
This approach however favors unbalanced partitions, leading to undesirable solutions. A way of addressing this is to consider the normalized cut value $\operatorname{\operatorname{NCut}}(A,B)$ [3] instead,
$\operatorname{\operatorname{NCut}}(A,B)=\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(A)}+\frac{\operatorname{\operatorname{Cut}}(A,B)}{\operatorname{\operatorname{Vol}}(B)},$
where $\operatorname{\operatorname{Vol}}(A)=\sum_{i\in A,j\in V}w(i,j)$ is the “volume” of $A$. We use $\operatorname{\operatorname{Vol}}(V)=\operatorname{\operatorname{Vol}}(A)+\operatorname{\operatorname{Vol}}(B)$ to represent the volume of $G$. Let $W$ be the weighted adjacency matrix of $G$ and $D$ be the diagonal degree matrix with $D(i,i)=\sum_{j\in V}w(i,j)$. The Laplacian matrix of $G$ is then defined as $L=D-W$. The problem of minimizing the $\operatorname{\operatorname{NCut}}$ of $G$ becomes:
$\min_{\mathbf{x}}$ $\mathcal{E}(\mathbf{x}|G),$ (1)
@@
2.2 Traditional and Recent Approaches for Edge Weighting and their Drawbacks
A crucial step in developing efficient segmentation solvers using $\operatorname{\operatorname{NCut}}$ is to define how the function $w$ is computed. In the original $\operatorname{\operatorname{NCut}}$ work and in early subsequent work [20, 21], the $w$ combined two grouping cues in a single value:
@@ 2.2 Traditional and Recent Approaches for Edge Weighting and their Drawbacks
where $I(j)$ and $X(j)$ are the appearance (such as the color or other pixel-level semantic feature) of pixel $j$ and its spatial location on image $I$, respectively, $\mathds{1}(\cdot)$ is the indicator function, and $r$, $\sigma_{I}$ and $\sigma_{X}$ have fixed values.
The $\operatorname{\operatorname{NCut}}$ formulation based on the above weight function presents several drawbacks. For one, it is computationally inefficient to construct the graph in Eq. 4 for high-resolution images, since it requires visiting each pixel and their neighboring locations within a radius $r$. Secondly, the graph is potentially dense, which poses memory and runtime issues during eigenvector computation. $\operatorname{\operatorname{NCut}}$ with weights as in Eq. 4 is also known to be underperforming in practice [14] and calibrating the values for $\sigma_{X}$ and $\sigma_{I}$ is challenging [18]. That led to the development of spectral image segmentation methods that (1) use superpixels [22] instead of working on the original pixel set for algorithmic efficiency and (2) employ contour-based information, such as Probability of Boundary (Pb) [14] instead of raw color intensities [15, 13] for performance. Finally, the optimization in Eq. 1 on the graph constructed using Eq. 4 lacks interpretability [18], i.e., all the understanding one can conceive out of its solutions is that it approximately finds segments whose pixels are similar in appearance and especially close enough to each other. Ideally, one would also be concerned about the extent to which certain parameters and each constituent data component (such as the pixel appearance data and its location information) affects the algorithm’s performance.
More recently, researchers took advantage of the performant but low-resolution features from self-supervised DINO architectures [16], to derive fully connected graphs whose weights are mostly given by simple cosine similarities [10, 11]. Although the resulting eigenvectors can be used to achieve good results after rescaling, such methodologies still rely on their low-resolution setting to be computationally efficient. Furthermore, this approach does not address the lack of interpretability discussed above.
Table 1: Results on $100\times 100$ synthetic images arising from different binary patterns. We use RGB features in each method. We outperform in runtime and mIoU classical spectral segmentation techniques in many challenging noise and pattern settings.
@@ 3.1 New Graph Formulation
Inspired by [2] and [4], we show that a sparse graph formulation efficiently addresses the issues above. Our graph construction starts by considering a grid graph $G_{\text{grid}}$ whose vertices correspond to image pixels, with edges of weight $\mu>0$ between each pair of neighboring pixels. From $G_{\text{grid}}$, we create a new graph $G_{\text{all}}$ by adding $k$ extra nodes, each corresponding to one of the $k$ unique colors in $I$ [^1] What is described here can be generalized to consider $k$ clusters of pixel-level features, instead of $k$ unique colors. This will be important when handling images where the number of unique colors is large or with multidimensional continuous pixel features, such as the ones arising from self-supervised vision transformers [16].. Each node on $G_{\text{grid}}$ is then connected to the extra node corresponding to its color with an edge of unitary weight. This graph formulation is computationally attractive for its sparsity. Assuming $k\ll n$, $G_{\text{all}}$ only requires $O(n)$ edges. When segmenting $I$, we only make use of the assignment vector for the grid nodes.
Image
Ours (RGB)
NCut
MixNCut
MGSC
Fig. 1: Sample qualitative results on synthetic images. Top: Gaussian noise, $\sigma_{g}^{2}=1.5$. Bottom: Salt & Pepper, $d=0.7$.
@@ 3.2 Interpretation of the Cut on $G_{\text{all}}$
Because of the grid nature of $G_{\text{grid}}$, the second term in Eq. 6 accounts for the number of neighboring pixel pairs that are assigned to different segments:
@@
Minimizing $\mathcal{E}(\mathbf{z}|G_{\text{all}})$ for $\mathbf{x}$ and $\mathbf{y}$, results in finding a partition of $G_{\text{all}}$ that trades off between spatial coherence (neighboring pixels should belong to the same segment) and color cohesion (pixels in a segment should have colors belonging to the same subset of available intensities) via the value of $\mu$. Minimizing the multiplicative factor in Eq. 9 also encourages a balanced partition of the nodes in $G_{\text{all}}$.
[…]
Fig. 2: Segmentation results for varying image sizes. Left: a sample image for $\ell=500$. Right: runtime and mIoU results varying the number of image pixels $n=\ell^{2}$.
Table 2: Average mIoU on BSDS500. We provide results for our method on RGB intensities or DINO pixel-level features. Our proposed method is shown to overperform many state-of-the-art unsupervised segmentation algorithms even in its simple RGB-based formulation. †The method listed was re-trained using code made available by their authors.
$\operatorname{\operatorname{NCut}}$ $K$-means GS IIC DFC† DoubleDIP† PiCIE† W-Net† SegSort† Ours (RGB) Ours (DINO)
0.385 0.240 0.313 0.172 0.305 0.356 0.325 0.428 0.480 0.479 0.534
Finally, one can spot some formulaic similarities between the expression in Eq. 6 and typical optimization objectives in inverse problem solvers [23], such as those involving the compromise of a quadratic data fidelity term and a regularization functional. Despite that similarity not being suitable for a complete interpretation of Eq. 6, since here we have both $\mathbf{x}$ and $\mathbf{y}$ as unknowns, it can still aid our understanding of the minimization carried out in our proposed graph. Minimizing Eq. 6 also corresponds to finding an assignment vector $\mathbf{z}$ where, on one hand, the extra node cluster assignments $\mathbf{y}$ are responsible as much as possible to the labeling of the pixel nodes $\mathbf{x}$ (first term), while also encouraging neighboring pixel labels to agree among themselves (second term).
@@ 4.1 Experimental Setup
We proceed with experiments on real and synthetic images. For our synthetic results, we contaminate three different background and foreground square patterns (intensities of either 0 and 1) of side length $\ell$ with either additive zero mean and variance $\sigma_{g}^{2}$ Gaussian noise or Salt & Pepper (S&P) with noise density $d$. The generated images are then rescaled and discretized, so the color intensities are integers in $[0,255]$. We also test our method using all images in BSDS500 [14], DUT-OMRON [24], DUTS [25] and ECSSD [26] segmentation datasets. We test segmenting them solely based on either their RGB data or their DINOv2 [9] features interpolated to the pixel level. In practice, we perform PCA on the extracted DINO features and the first 3 principal dimensions to be interpolated. The choice of 3 components is inspired by the original work in [9]. For each image, we vectorize their pixel feature space using mini-batch $K$-means clustering [27] with $k=256$ as the only preprocessing stage for our proposed method. Unless otherwise specified, we set $\mu=0.01$. In our methods, we do not apply any edge refinement post-processing.
Since the images in our real dataset may contain multiple objects/segments in their ground-truth, we generalized our proposed binary segmentation method taking the prescription in [3] as inspiration. Here, we recursively reapply our proposed method on the partitions found by it until the energy $\mathcal{E}(\mathbf{z}|G_{\text{all}})$ (ref. Eq. 5) of that partition is below given a threshold (set to 0.01) or until a given partition had already been subdivided 3 times.
Table 3: Average mIoU on binary segmentation datasets.
Method    DUTS    ECSSD DUT-OMRON
LOST 0.518 0.654 0.410
DeepUSPS 0.305 0.440 0.305
BigBiGAN 0.498 0.672 0.453
DeepSpectral 0.514 0.733 0.567
TokenCut 0.576 0.712 0.533
Ours (RGB) 0.577 0.617 0.509
Ours (DINO) 0.569 0.684 0.588
(a) Image
(b) $\mu=0.01$ ($0.30s$)
(c) $\mu=1$ ($0.17s$)
(d) $\mu=50$ ($0.91s$)
Fig. 3: Qualitative results on BSDS500 images for various $\mu$ using pixel RGB features. The proposed methods is performs fast image segmentation in various imaging settings despite its spectral nature. Its sole parameter ($\mu$) regulates the level of border detail in the final segmentation. (b)-(d) show the eigenvector found and (left) and the resp. segmentation (right). Both images are of size $481\times 321$ and the average runtime is shown in parentheses.
For performance evaluation, we follow the literature [10, 11, 19] and assess the segmentations via the intersection over union $\text{IoU}=|R\cap S|/|R\cup S|$, where $R$ and $S$ are a ground truth image region and its predicted segmentation, respectively. We compute the mean IoU (mIoU) by averaging the IoU values over the image [28]. The predicted segments are matched one-to-one with target segments using a version of the Hungarian algorithm modified to accommodate when the number of segments in the final segmentation is different from that of the ground truth. BSDS provides multiple ground-truth segmentations, so we average their respective mIoUs.
We compare our proposed method with $\operatorname{\operatorname{NCut}}$ and two more spectral methods for unsupervised segmentation, Multi Graph Spectral Clustering (MGSC) [29] and MixNCut [18] in our synthetic experiments. For our real data, we compare our results to those of three traditional baselines, $K$-means, Graph-based Segmentation (GS) [30] and SLIC [22]; and various deep learning-based unsupervised segmentation algorithms: W-Net [31], IIC [32], SegSort [33], DeepUSPS [34], DFC [28], LOST [35], PiCIE [27], BigBiGAN [36], DeepSpectral [10], TokenCut [11] and ACSeg [37]. When not otherwise specified as re-trained, we used the quantitative results provided in [28], [19] and [11]. W-Net, DeepSpectral and TokenCut are of particular interest to us, as they use Normalized Cuts concepts to devise their segmentation techniques.
Parameter tuning and model selection are performed in each re-trained method. We refrain from any sparsification algorithm in any spectral method. For $\operatorname{\operatorname{NCut}}$, we also set $r=\infty$ in Eq. 4 for simplicity. In our synthetic experiments, we apply $\operatorname{\operatorname{NCut}}$ on the full pixel set. In the experiments using RGB images, we first use SLIC [22] to over-segment the image in 500 superpixels. We also apply the recursive strategy above for $\operatorname{\operatorname{NCut}}$ when segmenting multi-region images.
@@
The experiments ran on an NVIDIA GeForce RTX 4070 GPU with 8 Gb VRAM memory and were implemented [^2] A demo code for this paper is available at www.github.com/rahul-palnitkar/Sparse-Graph-Spectral-Segmentation. in Python 3.10 using the GPU-accelerated sparse linear algebra utils of CuPy 12.3.